### v2

# Quantum Thermoelectric Circuits: A Universal Approach

Abstract
In this work, we develop a panoramic schematic of a quantum thermoelectric circuit theory in the steady state regime. We establish the foundations of the said premise by defining the analogs of Kirchhoff’s laws for heat currents and temperature gradients. We further show that our approach encompasses various circuits like thermal diode, transistor, and Wheatstone bridge. Additionally, we have been able to develop a model of a quantum thermal step transformer. We also construct a novel model of a thermal adder circuit, paving the way to develop thermal integrated circuits. This sheds new light on the present architecture of quantum device engineering.
Introduction.— All quantum systems are inherently subject to interactions with their surrounding environment, which the theory of open quantum systems addresses [1, 2, 3]. Recent technological advancements have elevated the importance of these interactions, particularly in quantum thermodynamics, where understanding non-equilibrium processes and energy flow at the quantum level has become essential for developing quantum technologies [4, 5, 6, 7, 8, 9].
Logical computational circuits are crucial for the development of quantum computers [10, 11, 12], making quantum circuitry foundational in quantum device engineering. Drawing from classical circuit design, identifying quantum equivalents of diodes, transistors, resistors, and inductors is vital for creating a quantum microprocessor. A promising approach involves using heat currents to create quantum versions of electrical and thermal devices [13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43]. Analyzing heat flow patterns in quantum networks is essential to model these quantum thermoelectric devices, focusing on identifying heat current behavior [44, 45]. From the advent of the formal theory of open quantum systems, an important achievement of which was the establishment of the Gorini-Kossakowski-Sudarshan-Lindblad (GKSL) master equation [46, 47], modeling quantum heat engines and thermal devices has been one of its most promising applications [48, 49], where heat currents are manipulated to efficient effect. This paper offers a novel quantum thermoelectric network theory, providing the fundamental laws on heat currents and temperature gradients, playing the equivalent roles of electric current and voltage differences, respectively.
We consider the thermoelectric quantum networks from the backdrop of Markovian dynamics [2, 46, 3]. The model involves multiple interacting qubits influenced by non-interacting thermal environmental modes, which can be represented as weak atom-photon couplings, with the baths considered as collections of harmonic oscillators [2]. In this backdrop, the resulting reduced dynamics of the system demonstrates a one-way information flow from the system to the environment, monotonically leading the qubit towards its corresponding thermal equilibrium. It is also important to mention that all our investigations are done in the steady state region, which naturally occurs for such Markovian dynamics. This, in turn, ensures the stability of the circuit models.
We model a quantum thermal resistor identified as the strength of interaction Hamiltonian to find the heat current between two or more qubits. We establish the laws of a thermoelectric circuit, analogous to Kirchhoff’s current and voltage laws, treating a qubit as the node. We use this framework to establish the balance condition of a quantum thermal Wheatstone bridge. In this discourse, we also prove the existence of a quantum thermal step transformer. We further develop a thermal adder circuit, motivated by the operational amplifier circuit, where the voltage output is the algebraic sum of all the inputs. It is also shown that the thermal diode and transistor circuits fall into the framework constructed here, the details of which are presented in the Supplemental Material [50]. The framework introduced here is the first of its kind, providing a fundamental tool to build a comprehensive quantum thermoelectric circuit theory.
Quantum thermal resistor and current law.— Our program begins by constructing a model for a quantum thermal resistor. By analyzing the flow of heat current through a quantum system between two thermal baths with differing temperatures, we seek to establish a relationship akin to Kirchhoff’s laws. Our goal is to understand if, and under what conditions, temperature gradients and heat currents in quantum thermal devices can relate in a manner analogous to these fundamental circuit principles. The quantum thermal resistor is made up of two qubits, Fig. 1(a).
Figure 1: Schematic diagrams of the (a) quantum thermal resistor and (b) 3-qubit quantum thermal circuit. Quantum thermal current from source $j$ to qubit $k$, taking qubit $k$ as a node, is $\mathcal{J}_{jk}$.
Qubits 1 and 2 are under the influence of bosonic baths at temperatures zero and $T$, respectively. The Hamiltonian of the two-qubit system (for $\hbar=k_{B}=1$) is given by $H_{S}=\sum_{j=1}^{2}\frac{\omega_{j}}{2}\sigma_{j}^{z}+J_{12}\left(\sigma^{x}_{1}\sigma^{x}_{2}+\sigma^{y}_{1}\sigma^{y}_{2}\right),$ where $\sigma^{i}~{}(i=x,y,z)$ are the Pauli matrices. $J_{12}$ is the interaction strength between the two qubits and models the quantum thermal resistor. The dynamics of the system is governed by the GKSL master equation of the form
$\frac{d\rho}{dt}$ $=-i[H_{S},\rho]+\mathcal{D}_{A1}\left(\rho\right)+\mathcal{D}_{B2}\left(\rho\right)=\mathcal{L}\left(\rho\right),$ (1)
where $\mathcal{D}_{jk}(\rho)=\gamma_{j}(\tilde{N}_{jk}+1)\left(\sigma^{-}_{k}\rho\sigma^{+}_{k}-\frac{1}{2}\left\{\sigma^{+}_{k}\sigma^{-}_{k},\rho\right\}\right)+\gamma_{j}\tilde{N}_{jk}\left(\sigma^{+}_{k}\rho\sigma^{-}_{k}-\frac{1}{2}\left\{\sigma^{-}_{k}\sigma^{+}_{k},\rho\right\}\right)$, with $\sigma^{\pm}_{k}=\frac{1}{2}\left(\sigma^{x}_{k}\pm i\sigma^{y}_{k}\right)$, and $\tilde{N}_{jk}=\frac{1}{e^{\beta_{j}\omega_{k}}-1}$, with $\beta_{j}=T_{j}^{-1}$ (here, $j=A,B$ and $k=1,2$). $\gamma_{A}$ and $\gamma_{B}$ are the dissipative factors. $\mathcal{L}$ denotes the right-hand side of the master equation. We assume that each of the qubits is locally interacting with its associated bath. The steady state of the system is given by the condition $\frac{d\rho^{SS}}{dt}=0$, and is
[…]
where the forms of $\alpha_{ij}$ and $\eta_{ij}$ are given in [50] and |$\ket{i}$ is in the computational basis. The quantum heat currents, in the steady state condition, from the bath $j$ to the qubit $k$ ($\mathcal{J}_{jk}$) and from qubit $l$ to qubit $k$ ($\mathcal{J}_{lk}$) are given by
$\mathcal{J}_{jk}={\rm Tr}\left[H_{k}\mathcal{D}_{jk}\left(\rho^{SS}\right)\right]~{}~{}\text{and}~{}~{}\mathcal{J}_{lk}=i{\rm Tr}\left\{\rho^{SS}[H_{lk},H_{k}]\right\},$ (3)
respectively [50].
Using the state $\rho^{SS}$ from Eq. (2), the current $\mathcal{J}_{A1}$ from bath $A$ to qubit 1 is $\mathcal{J}_{A1}={\rm Tr}\left[H_{1}\mathcal{D}_{A1}\left(\rho^{SS}\right)\right]=-\gamma_{A}\omega_{1}\left(\frac{\alpha_{00}}{\eta_{00}}+\frac{\alpha_{11}}{\eta_{11}}\right).$ Pertinently, since the parameters $\gamma_{A}\geq 0$, $\alpha_{00}/\eta_{00}\geq 0$ and $\alpha_{11}/\eta_{11}\geq 0$, the current $\mathcal{J}_{A1}$ negative. The rationale behind $\mathcal{J}_{A1}$ being negative is that qubit 1 is surrounded by a zero temperature bath, and negative $\mathcal{J}_{A1}$ indicates that the heat flow is directed from qubit 1 towards the thermal reservoir. Therefore, reservoir $A$ acts as a “ground” to which the quantum thermal resistor is connected. Furthermore, the current $\mathcal{J}_{21}$ from qubit 2 to 1 is given by $\mathcal{J}_{21}=\frac{2iJ_{12}\omega_{1}\left(\alpha_{12}-\alpha_{12}^{*}\right)}{\eta_{12}}=\frac{-4J_{12}\omega_{1}\Im\left(\alpha_{12}\right)}{\eta_{12}},$ where $\Im(*)$ denotes imaginary part. Interestingly, we find $\mathcal{J}_{A1}=-\left(\frac{\gamma_{A}\omega_{1}x_{0}}{\gamma_{B}}\right)\frac{\alpha_{00}}{\eta_{00}}=-\mathcal{J}_{21}$, resulting in $\mathcal{J}_{A1}+\mathcal{J}_{21}=0.$ This is crucial as it allows qubit 1 to be considered as a junction node, leading to the visualization of the above equation as the quantum thermal version of Kirchhoff’s current law. In fact, considering heat currents $\mathcal{J}_{jk}$’s from $n$ different sources to a qubit $k$, it can be proved that in the steady state condition (see [50]), the sum of all the heat currents is zero, i.e.,
$\sum_{j=1}^{n}\mathcal{J}_{jk}=0.$ (4)
In a similar manner, we find that the currents with qubit 2 as the junction node are given by $\mathcal{J}_{B2}={\rm Tr}\left[H_{2}\mathcal{D}_{B2}\left(\rho^{SS}\right)\right]=\left(\frac{\gamma_{A}\omega_{2}x_{0}}{\gamma_{B}}\right)\frac{\alpha_{00}}{\eta_{00}}$, and $\mathcal{J}_{12}=-\left(\frac{\gamma_{A}\omega_{2}x_{0}}{\gamma_{B}}\right)\frac{\alpha_{00}}{\eta_{00}}.$ It can be easily verified that $\mathcal{J}_{B2}+\mathcal{J}_{12}=0$, which, upon assuming qubit 2 as a junction node, shows the characteristics of the quantum thermal version of Kirchoff’s current law.
Quantum thermal transformer.— Interestingly, we observe that the heat currents $\mathcal{J}_{21}$ and $\mathcal{J}_{12}$ are interconnected by the relation $\frac{\mathcal{J}_{21}}{\omega_{1}}=-\frac{\mathcal{J}_{12}}{\omega_{2}},$ or $\left|\frac{\mathcal{J}_{21}}{\mathcal{J}_{12}}\right|=\frac{\omega_{1}}{\omega_{2}}=\left|\frac{\mathcal{J}_{A1}}{\mathcal{J}_{B2}}\right|.$ In general, for an interaction $H_{jk}=J_{jk}\left(\sigma^{x}_{j}\sigma^{x}_{k}+\sigma^{y}_{j}\sigma^{y}_{k}\right)$ between two qubits $j$ and $k$, the current between them can be shown to be related by
$\frac{\mathcal{J}_{jk}}{\omega_{k}}=-\frac{\mathcal{J}_{kj}}{\omega_{j}}.$ (5)
The proof is given in [50]. The above relation mimics the quantum thermal version of a step transformer from electrical circuits, where the ratio of electric currents between two coils depends on the number of turns in each coil. Here, the ratio of quantum heat currents depends on the transition frequencies $\omega_{j}$ and $\omega_{k}$ of the qubits $j$ and $k$, respectively, thereby allowing for the manipulation of quantum heat currents.
Quantum thermal voltage.— We now develop the quantum analog of the thermal voltage driving the thermal current. To this end, we determine the effective temperature of the qubits 1 and 2 in the steady state. The reduced state of qubit 1 from the steady state, Eq. 2, is given by $\rho^{SS}_{1}=\left(\rho^{SS}_{00}+\rho^{SS}_{11}\right)\ket{0}\bra{0}+\left(\rho^{SS}_{22}+\rho^{SS}_{33}\right)\ket{1}\bra{1}.$ | . The effective temperature $T_{1}$ of the qubit 1 is found by comparing the state $\rho^{SS}_{1}$ with the state $e^{-H_{1}/T_{1}}/{\rm Tr}\left(e^{-H_{1}/T_{1}}\right)$, where $H_{1}=\frac{\omega_{1}}{2}\sigma^{z}_{1}$, leading to $T_{1}=\omega_{1}/\log\left(\frac{1}{\rho^{SS}_{00}+\rho^{SS}_{11}}-1\right)=\omega_{1}/\log\left(\frac{\gamma_{B}\eta_{00}}{x_{0}\alpha_{00}}-1\right).$ Note that $T_{1}$ can be envisaged as the thermal potential difference between the qubit 1 and bath $A$, that is, $V_{1A}=T_{1}-0=T_{1}$. This potential difference drives the thermal current $\mathcal{J}_{A1}$, which explicitly in terms of $T_{1}$ can be written as $\mathcal{J}_{A1}=\frac{-\gamma_{A}\omega_{1}}{1+e^{\omega_{1}/T_{1}}}.$ In the low and high-temperature limits, this is
$\mathcal{J}_{A1}=\begin{cases}\frac{-\gamma_{A}\omega_{1}}{2},&\text{for}~{}~{}T_{1}\gg 0\\ -\gamma_{A}\omega_{1}e^{-\omega_{1}/T_{1}}.&\text{for}~{}~{}T_{1}\sim 0\end{cases}$ (6)
Figure 2: Variation of the thermal current $|\mathcal{J}_{A1}|$ between the qubit 1 and the bath $A$ as a function of the thermal potential difference $V_{1A}=T_{1}$ for $\gamma_{A}=0.5$ and $\omega_{1}=1.0$.
This has been depicted in Fig. 2. Similarly, one can find the reduced state of the qubit 2 and its corresponding effective temperature, which is $T_{2}=\omega_{2}/\log\left(\frac{1}{\rho^{SS}_{00}+\rho^{SS}_{22}}-1\right).$ The potential difference between the bath $B$ and the qubit 2 can now be written as $V_{B2}=T-T_{2}$. Further, $\mathcal{J}_{B2}=\frac{\gamma_{B}\omega_{2}}{2}\left[-1+\coth\left(\frac{\omega_{2}}{2T}\right)\tanh\left(\frac{\omega_{2}}{2\left\{T-V_{B2}\right\}}\right)\right].$ It can be pointed out here that the thermal current $\mathcal{J}_{B2}$ does not explicitly depend only on the thermal potential difference $V_{B2}$ but also on the temperature ($T$) of the bath $B$.
Figure 3: Variation of heat currents $\mathcal{J}_{jk}$ in (a), and heat current $\mathcal{J}_{21}$ from qubit 2 to 1 in (b), (c), and (d) with temperature $T$ of bath $B$ for the two-qubit quantum thermal resistor model. (a) $\omega_{1}=2,\omega_{2}=2.5,\gamma_{A}=0.01,\gamma_{B}=0.05$, and $J_{12}=1.0$; (b) $\omega_{1}=5,\gamma_{A}=0.01,\gamma_{B}=0.005$, and $J_{12}=1.0$; (c) $\omega_{1}=2,\omega_{2}=5,\gamma_{A}=0.01$, and $J_{12}=1.0$; (d) $\omega_{1}=1.0,\omega_{2}=15$, and $\gamma_{A}=\gamma_{B}=0.01$.
The heat currents $\mathcal{J}_{jk}$ discussed here are plotted in Fig. 3 as a function of temperature $T$. The heat current can be seen to increase linearly in certain parameter regimes but, in general, exhibits non-linear behavior. However, Kirchhoff’s current law is satisfied universally.
Law of thermal potentials.— Having discussed quantum thermal resistors, we now move on to study quantum thermal circuits made up of multiple qubits. For this purpose, we consider quantum thermal circuits composed of three qubits. A schematic diagram of this circuit is given in Fig.  1(b), where qubits 1 and 2 are under the influence of bosonic baths $A$ and $B$ at zero and $T$ temperatures, respectively. In this setup, we consider a third qubit that interacts with qubits 1 and 2, which themselves interact with each other. The Hamiltonian for this three-qubit setup is $\widetilde{H}_{S}=\frac{1}{2}\sum_{i=1}^{3}\omega_{i}\sigma^{z}_{i}+\sum_{l,k=1,l<k}^{3}J_{lk}\left(\sigma^{x}_{l}\sigma^{x}_{k}+\sigma^{y}_{l}\sigma^{y}_{k}\right),$ where qubits $l$ and $k$ interact with each other via Heisenberg $XX$ type interaction with the interaction strength $J_{lk}$ and $\omega_{k}$ being the transition frequency of the $k$-th qubit. We compute the steady-state $\tilde{\rho}^{SS}$ of this system using Eq. (1) by replacing $H_{S}$ with $\widetilde{H}_{S}$.
Figure 4: Variation of quantum heat currents in the case of the three-qubit quantum thermal circuit in steady state. The parameters are: $\omega_{1}=1.25,\omega_{2}=1.5,\omega_{3}=1.75,J_{12}=1.0,J_{13}=0.5,J_{23}=0.75,\gamma_{A}=0.1,\gamma_{B}=0.05$.
Figure 5: Schematic diagrams of the (a) quantum thermal Wheatstone Bridge and (b) quantum thermal adder circuit.
The variation of the quantum heat currents in this circuit is shown in Fig. 4. It can be verified from this figure that Kirchhoff’s current law, which is discussed above, holds true for quantum thermal currents at a junction node. Further, the effective temperatures $T_{1}$, $T_{2}$, and $T_{3}$ of the qubits 1, 2, and 3, respectively, can be calculated using the steady-state $\tilde{\rho}^{SS}$ of the three-qubit system similarly to the previous case, and are given by $T_{1}=\omega_{1}/\log\left(\frac{1}{\tilde{\rho}^{SS}_{00}+\tilde{\rho}^{SS}_{11}+\tilde{\rho}^{SS}_{22}+\tilde{\rho}^{SS}_{33}}-1\right)$, $T_{2}=\omega_{2}/\log\left(\frac{1}{\tilde{\rho}^{SS}_{00}+\tilde{\rho}^{SS}_{11}+\tilde{\rho}^{SS}_{44}+\tilde{\rho}^{SS}_{55}}-1\right)$, and $T_{3}=\omega_{3}/\log\left(\frac{1}{\tilde{\rho}^{SS}_{00}+\tilde{\rho}^{SS}_{22}+\tilde{\rho}^{SS}_{44}+\tilde{\rho}^{SS}_{66}}-1\right)$. The corresponding thermal potential differences are given by $V_{1A}=T_{1}-0;V_{21}=T_{2}-T_{1};V_{32}=T_{3}-T_{2};V_{13}=T_{1}-T_{3};V_{B2}=T-T_{2}.$ Using this, we get a relation for the potential differences inside the loop of the three qubits, which is $V_{21}+V_{32}+V_{13}=0$, resembling Kirchhoff’s second law of an electric circuit, which states that the sum of potential differences around a closed loop in an electric circuit is zero. The relation between the thermal current $\mathcal{J}_{A1},\mathcal{J}_{B2}$ and the potential difference $V_{1A},V_{B2}$ is the same as in the case of the previous circuit. In terms of the elements of the steady state $\tilde{\rho}^{SS}$ of the three-qubit circuit, the quantum thermal currents from qubit 3 to 1 and from qubit 2 to 1 are given by $\mathcal{J}_{31}=-4J_{13}\omega_{1}\left[\Im{(\tilde{\rho}^{SS}_{14})}+\Im{(\tilde{\rho}^{SS}_{36})}\right]$ and $\mathcal{J}_{21}=-4J_{12}\omega_{1}\left[\Im{(\tilde{\rho}^{SS}_{24})}+\Im{(\tilde{\rho}^{SS}_{35})}\right]$, respectively. This leads to a relation between the quantum thermal heat current $\mathcal{J}_{31}$ ($\mathcal{J}_{13}$) and the thermal potential $V_{13}$ given by
$\mathcal{J}_{31}=\frac{4J_{13}}{\tilde{f}_{13}}V_{13}=-\frac{\omega_{1}\mathcal{J}_{13}}{\omega_{3}},$ (7)
where the $\tilde{f}_{13}$ is a function of $\omega_{j},T_{j},J_{ij},T$, and $\gamma_{k}$ (for $i=1,2;j=1,2,3$ and $k=A,B$). Details of the proof are presented in [50]. Similarly, one can find out the various functions, such as $\tilde{f}_{12}$, and $\tilde{f}_{23}$, which can be used to identify the corresponding relation of the quantum thermal currents $\mathcal{J}_{12},\mathcal{J}_{21},\mathcal{J}_{23}$, and $\mathcal{J}_{32}$ with the quantum thermal potentials $V_{12}$ and $V_{23}$.
We use the above framework to develop the quantum thermal equivalent of an electric Wheatstone bridge and a quantum thermal adder circuit. Also, a diode and a quantum thermal transistor, from this perspective, are discussed in [50].
Quantum thermal Wheatstone bridge.— The classical Wheatstone bridge is a device primarily used to determine an unknown resistance. In a circuit of four resistances, one is a tunable resistance, and the other is the unknown, whose value is determined when the ratio between them becomes equal to that of the known resistances. In the case of the quantum Wheatstone bridge, the proposed device would determine an unknown interaction strength of a Hamiltonian. A pictorial comparison between a classical Wheatstone bridge and a quantum one is depicted in [51], wherein a model of quantum Wheatstone bridge was proposed, with the balance condition determining the unknown coupling strength obtained by observing a drop in population of Bell states caused by controlling a tunable coupling. On the other hand, our approach to finding the balance condition is via manipulating the heat current through the circuit, in tune with the actual classical Wheatstone bridge.
The Hamiltonian for the quantum Wheatstone bridge is given by $H_{S}^{W}=\frac{1}{2}\sum_{i=1}^{4}\omega_{i}\sigma^{z}_{i}+\sum_{l,k=1,l<k}^{4}J_{lk}\left(\sigma^{x}_{l}\sigma^{x}_{k}+\sigma^{y}_{l}\sigma^{y}_{k}\right)$, where $J_{12}=0$ as the qubits 1 and 2 are non-interacting. A schematic diagram of the quantum thermal Wheatstone bridge is shown in Fig. 5(a). This is similar to the setup of a Wheatstone bridge of an electrical circuit, where an unknown resistance can be determined using a specific balance condition when the current between nodes 3 and 4 is zero. Qubits 1 and 2 are impacted by the baths $A$ at zero temperature and $B$ at temperature $T,$ respectively. The evolution of the system is dictated by the quantum master equation of the form given in Eq. 1 by replacing $H_{S}$ with $H_{S}^{W}$, where the steady-state $\rho^{SS}_{W}$ can be obtained using $\mathcal{L}\left(\rho^{SS}_{W}\right)=0$. The balance condition for the Wheatstone bridge is that the quantum heat current $\mathcal{J}_{34}$ between qubits 3 and 4 should be zero. This happens when couplings follow the relation: $J_{13}=J_{14}$ and $J_{23}=J_{24}$. Note that $J_{13}$ need not be equal to $J_{23}$.
Figure 6: Variation of quantum heat current $\mathcal{J}_{43}$ between qubits 3 and 4 in the case of the quantum Wheatstone bridge model in the steady state. The green diamonds show the condition derived in [51], and the red circles show the balance condition discussed in this paper. The parameters are kept the same as in [51] for comparison: $\omega_{1}=\omega+2h_{1},\omega_{2}=\omega_{3}=\omega,\omega_{4}=\omega+2h_{4}$; $\omega=20,J_{34}=20J,h_{1}=20J,h_{4}=0.5J,\gamma_{A}=J,\gamma_{B}=10J,J_{23}=J_{24}=J,J_{13}=2J$; and $J=0.1$, where $h_{i}$’s are an offset in the magnetic field acting on the $i$-th spin.
The phenomenon of the quantum Wheatstone Bridge is shown in Fig.  6. It can be observed that for the condition $J_{13}=J_{14}$, the red markers, depicting the current between qubits 3 and 4, are always zero. Further, green markers show the balance condition derived in [51], which is consistent at low temperatures; however, at high temperatures, the green markers shift slightly above the zero line, deviating from the balance condition.
Quantum thermal adder circuit.– We now move on to discuss the quantum adder circuit, Fig. 5(b). This thermal circuit is motivated by the DC operational amplifier’s adder circuit, where the output voltage of the device is the sum of all the input voltages. Here, the qubits $1,2,...,n$ interact with the baths $I,II,...,N$, at temperatures $T_{I},T_{II},...,T_{N}$, respectively. The effective temperature of the qubits $1,2,...,n$ are $T_{1},T_{2},...T_{n}$, respectively. The thermal potential difference between the qubit 1 and bath $I$ is $V_{I1}=T_{I}-T_{1}$ and similarly for the other qubits. The qubits $1,2,...,n$ also interact with the qubit $\alpha$, which is connected to a zero temperature bath $T_{A}$ and the thermal potential difference between qubit $\alpha$ and bath $A$ is $V_{\alpha A}=T_{\alpha}-T_{A}=T_{\alpha}$, where $T_{\alpha}$ is the effective temperature of qubit $\alpha$. The Hamiltonian of the system is given by $H_{add}=\frac{\omega_{\alpha}}{2}\sigma^{z}_{\alpha}+\sum_{i=1}^{n}\left[\frac{\omega_{i}}{2}\sigma^{z}_{i}+J_{i\alpha}\left(\sigma^{x}_{\alpha}\sigma^{x}_{i}+\sigma^{y}_{\alpha}\sigma^{y}_{i}\right)\right]$. From Kirchhoff’s thermal current law, taking qubit $\alpha$ as a node, we can write $\mathcal{J}_{A\alpha}=\sum_{k=1}^{n}\mathcal{J}_{k\alpha}$ and using the relation $\mathcal{J}_{\alpha k}/\omega_{k}=-\mathcal{J}_{k\alpha}/\omega_{\alpha}$ (for $k=1,2,..,n$) from quantum thermal transformer, we can write $\mathcal{J}_{A\alpha}=-\sum_{k=1}^{n}\frac{\omega_{\alpha}}{\omega_{k}}\mathcal{J}_{\alpha k}$. But $\mathcal{J}_{\alpha k}=\mathcal{J}_{jk}$ (for $(j,k)=(I,1),(II,2),...(N,n)$); therefore, we get a relation between the input thermal currents $\mathcal{J}_{jk}$ and the output thermal current $\mathcal{J}_{A\alpha}$ as
$\mathcal{J}_{A\alpha}=-\omega_{\alpha}\left(\frac{\mathcal{J}_{I1}}{\omega_{1}}+\frac{\mathcal{J}_{II2}}{\omega_{2}}+...+\frac{\mathcal{J}_{Nn}}{\omega_{n}}\right),$ (8)
which for $\omega_{k}=\omega_{\alpha}\forall k$ results into $\mathcal{J}_{A\alpha}=-\left(\mathcal{J}_{I1}+\mathcal{J}_{II2}+...+\mathcal{J}_{Nn}\right)$. The above equation exactly mimics the relation between the input voltage and the output voltage of an adder circuit. As seen before, the currents $\mathcal{J}_{jk}$ and $\mathcal{J}_{A\alpha}$ are related to the effective temperature of the corresponding qubits $T_{k}$ via $\mathcal{J}_{jk}=\gamma_{j}\omega_{k}\left[-1+\coth\left(\frac{\omega_{k}}{2T_{j}}\right)\tanh\left(\frac{\omega_{1}}{2T_{k}}\right)\right],$ for $j=A,I,II,...N$ and $k=\alpha,1,2,...n$. Taking $\gamma_{A}=\gamma_{I}=...=\gamma_{N}$, $\omega_{1}=\omega_{2}=...=\omega_{n}=\omega$, $J_{1\alpha}=J_{2\alpha}=...=J_{n\alpha}$ and $T_{I}=T_{II}=...=T_{N}=T$, we get a simplified relation between the effective temperatures ($T_{x}$ for all qubits) of the qubits $1,2,...,n$ and the qubit $\alpha$ ($T_{\alpha}$), i.e.,
$\tanh\left(\frac{\omega}{2T}\right)\left(N+\frac{2}{1+e^{\omega_{\alpha}/T_{\alpha}}}\right)=N\tanh\left(\frac{\omega}{2T_{x}}\right).$ (9)
Immediately, we observe that at lower temperatures $T$, the factor $\frac{2}{1+e^{\omega_{\alpha}/T_{\alpha}}}\rightarrow 0$. Interestingly, for this scenario, for higher values of $T$ and lower values of $\omega_{\alpha}$, the potential difference between bath $\mathcal{B}$ and qubits $V_{\mathcal{B}x}=T-T_{x}$ becomes independent of $T_{\alpha}$ and reduces to $V_{\mathcal{B}x}\approx\frac{T}{N+1}$. Remarkably, under certain parameter regimes, we get the condition that the sum of input thermal potentials becomes approximately equal to the output potential. For example, in the case of two qubits 1 and 2 interacting with qubit $\alpha$, keeping $\omega_{1}=\omega_{2}=\omega$, $J_{1\alpha}=J_{2\alpha}$, $T_{I}=T_{II}=T$, and $\gamma_{I}=\gamma_{II}=\gamma$, we get a condition on $\omega_{\alpha}$ ($\omega_{\alpha}=1.618N\omega$), and $\gamma_{A}$ ($\gamma_{A}=N\gamma$), such that $V_{I1}+V_{II2}=V_{\alpha A}$ with high accuracy in the high-temperature ($T$) regime. Similarly, for a higher number of qubits, $\omega_{\alpha}$ and $\gamma_{\alpha}$ can be tuned appropriately to get $V_{I1}+V_{II2}+...+V_{Nn}=V_{\alpha A}$, in general.
Conclusion.— In this paper, we have presented a comprehensive schematic for a quantum thermal circuit theory, paving a possible roadmap for the construction of quantum thermal integrated circuits, important ingredients of which are quantum thermal diodes and transistors. We prove the existence of Kirchhoff-like voltage and current laws, with interaction strength playing the role of circuit resistance. We further demonstrate how heat currents can be manipulated in these quantum circuits to construct a quantum step transformer. Continuing this line of study, we demonstrate a Quantum Wheatstone bridge, where unknown Hamiltonian strengths can be determined by the profiling of heat currents, which could have a plethora of applications in process tomography, metrology, and quantum sensing. A model of a quantum thermal adder circuit for temperature gradients was also constructed, creating further opportunities to design thermal operational amplifiers and other integrated circuits. The framework of the quantum thermoelectric circuit developed here can be adapted to experimental realization. Thus, for example, progress has been made towards the diode architecture with pentamethyl-disilane in an NMR register [52]. Along similar lines, experimental demonstration of other quantum thermoelectric circuits discussed here could be envisaged. It is thus sufficient to say that the present work offers the fundamentals of a comprehensive quantum thermoelectric circuit theory, paving the way for the experimental demonstration of quantum thermoelectric integrated circuits.
Acknowledgements.— SB2 acknowledges the support from the Ministry of Electronics and Information Technology, Government of India, under Grant No. 4(3)/2024-ITEA. SB1 acknowledges useful discussions with T.S. Mahesh on the possible experimental realizations of quantum thermal devices.
Supplemental Material for “Quantum Thermoelectric Circuits: A Universal Approach”
I Steady-state of quantum thermal resistor
In the following, we discuss the quantum thermal resistor. The Hamiltonian of the system made up of two qubits (for $\hbar=k_{B}=1$) is given by
$H_{S}$ $=H_{1}+H_{2}+H_{12}=\frac{\omega_{1}}{2}\sigma_{1}^{z}+\frac{\omega_{2}}{2}\sigma_{2}^{z}+J_{12}\left(\sigma^{x}_{1}\sigma^{x}_{2}+\sigma^{y}_{1}\sigma^{y}_{2}\right),$ (10)
where $\omega_{1}$ and $\omega_{2}$ are the transition frequencies for qubits 1 and 2, respectively, and $\sigma^{k}$’s (for $k=x,y,z$) are the Pauli spin matrices. $J_{12}$ is the interaction strength between qubits 1 and 2. To discuss the thermal transport between the two qubits and their respective baths, we find out the steady state of the system. The dynamics of the system under the Born-Markov and the secular approximations are given by the Gorini-Kossakowski-Sudarshan-Lindblad (GKSL) master equation of the form.
$\frac{d\rho}{dt}$ $=-i[H_{S},\rho]+\mathcal{D}_{A1}\left(\rho\right)+\mathcal{D}_{B2}\left(\rho\right),$
$=-i[H_{S},\rho]+\gamma_{A}\left(\sigma^{-}_{1}\rho\sigma^{+}_{1}-\frac{1}{2}\left\{\sigma^{+}_{1}\sigma^{-}_{1},\rho\right\}\right)$
$+\gamma_{B}\left(N_{th,B2}+1\right)\left(\sigma_{2}^{-}\rho\sigma_{2}^{+}-\frac{1}{2}\left\{\sigma^{+}_{2}\sigma^{-}_{2},\rho\right\}\right)$
$+\gamma_{B}N_{th,B2}\left(\sigma^{+}_{2}\rho\sigma^{-}_{2}-\frac{1}{2}\left\{\sigma^{-}_{2}\sigma^{+}_{2},\rho\right\}\right),$ (11)
where $\sigma^{\pm}_{j}=\frac{1}{2}\left(\sigma^{x}_{j}\pm i\sigma^{y}_{j}\right)$, and $N_{th,B2}=\frac{1}{e^{\omega_{2}/T_{B}}-1}$ with $T_{B}=T$. $\gamma_{A}$ and $\gamma_{B}$ are the dissipative factors. The steady state of the system is given by the condition $\frac{d\rho^{SS}}{dt}=0$. Here, baths $A$ and $B$ are at zero and $T$ temperatures, respectively. The form of the density matrix $\rho^{SS}$ that satisfies the said condition is
$\rho^{SS}=\begin{pmatrix}\frac{\alpha_{00}}{\eta_{00}}&&0&&0&&0\\ 0&&\frac{\alpha_{11}}{\eta_{11}}&&\frac{\alpha_{12}}{\eta_{12}}&&0\\ 0&&\frac{\alpha_{12^{*}}}{\eta_{12}}&&\frac{\alpha_{22}}{\eta_{22}}&&0\\ 0&&0&&0&&1-\frac{1}{1+e^{\beta\omega_{2}}}+\frac{\alpha_{33}}{\eta_{33}}\end{pmatrix},$ (12)
where
$\alpha_{00}$ $=16J_{12}^{2}\gamma_{B}^{2}\left[\gamma_{A}+\gamma_{B}\coth\left(\frac{\beta\omega_{2}}{2}\right)\right],$
$\eta_{00}$ $=x_{0}\left[\gamma_{A}\left\{16J_{12}^{2}x_{0}+\gamma_{B}\left(1+e^{\beta\omega_{2}}\right)\left(\gamma_{A}^{2}+4(\omega_{1}-\omega_{2})^{2}\right)\right\}\right.+\left.\gamma_{B}\coth\left(\frac{\beta\omega_{2}}{2}\right)\left\{16J_{12}^{2}x_{0}+\gamma_{A}\gamma_{B}\left(1+e^{\beta\omega_{2}}\right)\left(2\gamma_{A}+\gamma_{B}\coth\left(\frac{\beta\omega_{2}}{2}\right)\right)\right\}\right],$
$\alpha_{11}$ $=\left(x_{0}-\gamma_{B}\right)\alpha_{00},~{}~{}~{}\eta_{11}=\gamma_{B}\eta_{00},~{}~{}~{}\alpha_{12}=-4J_{12}\gamma_{A}\gamma_{B}\left[i\left\{\gamma_{A}+\gamma_{B}\coth\left(\frac{\beta\omega_{2}}{2}\right)\right\}+2\left(\omega_{1}-\omega_{2}\right)\right],$
$\eta_{12}$ $=\frac{\eta_{00}}{x_{0}},$
$\alpha_{22}$ $=\gamma_{B}\left[16J_{12}^{2}\left(e^{\beta\omega_{2}}-1\right)\left(x_{0}-\gamma_{B}\right)+\gamma_{A}\left\{x_{0}^{2}+4\left(e^{\beta\omega_{2}}-1\right)^{2}\left(\omega_{1}-\omega_{2}\right)^{2}\right\}\right],$
$\eta_{22}$ $=\frac{\eta_{00}\left(e^{\beta\omega_{2}}-1\right)^{2}}{x_{0}},$
$\alpha_{33}$ $=16J_{12}^{2}\left[\gamma_{A}+\gamma_{B}\coth\left(\frac{\beta\omega_{2}}{2}\right)\right]\left[\gamma_{A}^{2}\left(e^{\beta\omega_{2}}-1\right)^{2}-e^{\beta\omega_{2}}\left(e^{\beta\omega_{2}}+1\right)\gamma_{B}^{2}\right],$
$\eta_{33}$ $=\eta_{00}\left(e^{\beta\omega_{2}}+1\right),$ (13)
with $x_{0}=\gamma_{B}-\gamma_{A}+e^{\beta\omega_{2}}\left(\gamma_{A}+\gamma_{B}\right)$.
II Details on the forms of heat currents and proof of quantum thermal version of Kirchhoff’s current law
Consider a general system of $n$ qubits (labeled $1$, $2$, $3$, …, $n$), where each qubit is coupled to all other qubits. Further, each qubit is weakly coupled to its respective bosonic thermal bath (labeled $I$, $II$, $III$, …, $N$). The system’s Hamiltonian in this setup is given by
$H_{S}=\sum_{k=1}^{n}H_{k}+\sum_{l,k=1,l<k}^{n}H_{lk},$ (14)
where $H_{k}=\frac{\omega_{k}}{2}\sigma^{z}_{k}$ and $H_{lk}=J_{lk}\left(\sigma^{x}_{l}\sigma^{x}_{k}+\sigma^{y}_{l}\sigma^{y}_{k}\right)$ is the qubit-qubit interaction Hamiltonian with $J_{lk}$ being the coupling strength. Under the Born-Markov and rotating wave approximations, the dynamics of the system (depicted by $\rho$) is given by the GKSL master equation
$\frac{d\rho}{dt}=-i\left[H_{S},\rho\right]+\mathcal{D}_{I1}(\rho)+\mathcal{D}_{II2}(\rho)+...+\mathcal{D}_{Nn}(\rho),$ (15)
where $\mathcal{D}_{jk}(\rho)=\gamma_{j}(\tilde{N}_{jk}+1)\left(\sigma^{-}_{k}\rho\sigma^{+}_{k}-\frac{1}{2}\left\{\sigma^{+}_{k}\sigma^{-}_{k},\rho\right\}\right)+\gamma_{j}\tilde{N}_{jk}\left(\sigma^{+}_{k}\rho\sigma^{-}_{k}-\frac{1}{2}\left\{\sigma^{-}_{k}\sigma^{+}_{k},\rho\right\}\right)$, and $\tilde{N}_{jk}=\frac{1}{e^{\beta_{j}\omega_{k}}-1}$, with $\beta_{j}=T_{j}^{-1}$ for $(j,k)=(I,1),(II,2),...,(N,n)$. Let us pick a qubit $m$ (Hamiltonian $H_{m}$) from the $n$ qubits and consider it to be a node. This qubit $m$ would have quantum heat currents from all the sources, that is, from all the other qubits that interact with it as well as from the bosonic bath it is in interaction with (let the corresponding bath be $M$). The time derivative of the expectation value of $H_{m}$ is
$\frac{d}{dt}{\rm Tr}\left[H_{m}\rho\right]={\rm Tr}\left[\frac{dH_{m}}{dt}\rho\right]+{\rm Tr}\left[H_{m}\frac{d\rho}{dt}\right],$ (16)
where the left-hand side of the above equation denotes the rate of change in the energy of the qubit $m$. On the right-hand side (RHS), the first term is the power, and the second term denotes the net heat current directed towards qubit $m$. Note that, in the present scenario, the Hamiltonian $H_{m}$ is constant; therefore, the first term on the RHS of the above equation becomes zero. Thus, $\frac{d}{dt}{\rm Tr}\left[H_{m}\rho\right]={\rm Tr}\left[H_{m}\frac{d\rho}{dt}\right]$. Now, we multiply Eq. (15) with $H_{m}$ as
$H_{m}\frac{d\rho}{dt}$ $=-iH_{m}\left[H_{S},\rho\right]+H_{m}\mathcal{D}_{I1}(\rho)+H_{m}\mathcal{D}_{II2}(\rho)$
$+...+H_{m}\mathcal{D}_{Nn}(\rho).$ (17)
Taking trace on both sides in the above equation, we get
${\rm Tr}\left[H_{m}\frac{d\rho}{dt}\right]$ $=-i{\rm Tr}\left[H_{m}H_{S}\rho-H_{m}\rho H_{S}\right]+{\rm Tr}\left[H_{m}\mathcal{D}_{I1}(\rho)\right]$
$+{\rm Tr}\left[H_{m}\mathcal{D}_{II2}(\rho)\right]+...+{\rm Tr}\left[H_{m}\mathcal{D}_{Nn}(\rho)\right].$ (18)
The trace ${\rm Tr}\left[H_{m}\mathcal{D}_{jk}(\rho)\right]$ for $(j,k)=(I,1),(II,2),...,(N,n)$ in the above equation can be expanded as
${\rm Tr}\left[H_{m}\mathcal{D}_{jk}(\rho)\right]$ $={\rm Tr}\left[\gamma_{j}(\tilde{N}_{jk}+1)H_{m}\left(\sigma^{-}_{k}\rho\sigma^{+}_{k}-\frac{1}{2}\left\{\sigma^{+}_{k}\sigma^{-}_{k},\rho\right\}\right)\right]$
$+{\rm Tr}\left[\gamma_{j}\tilde{N}_{jk}H_{m}\left(\sigma^{+}_{k}\rho\sigma^{-}_{k}-\frac{1}{2}\left\{\sigma^{-}_{k}\sigma^{+}_{k},\rho\right\}\right)\right],$ (19)
which upon substituting $H_{m}=\frac{\omega_{m}}{2}\sigma^{z}_{m}$ becomes
${\rm Tr}\left[H_{m}\mathcal{D}_{jk}(\rho)\right]$ $=\gamma_{j}(\tilde{N}_{jk}+1)\frac{\omega_{m}}{2}{\rm Tr}\left[\left(\sigma^{z}_{m}\sigma^{-}_{k}\rho\sigma^{+}_{k}\right.\right.$
$-\left.\left.\frac{1}{2}\left\{\sigma^{z}_{m}\sigma^{+}_{k}\sigma^{-}_{k}\rho+\sigma^{z}_{m}\rho\sigma^{+}_{k}\sigma^{-}_{k}\right\}\right)\right]$
$+\gamma_{j}\tilde{N}_{jk}\frac{\omega_{m}}{2}{\rm Tr}\left[\left(\sigma^{z}_{m}\sigma^{+}_{k}\rho\sigma^{-}_{k}\right.\right.$
$-\left.\left.\frac{1}{2}\left\{\sigma^{z}_{m}\sigma^{-}_{k}\sigma^{+}_{k}\rho+\sigma^{z}_{m}\rho\sigma^{-}_{k}\sigma^{+}_{k}\right\}\right)\right].$ (20)
It can be easily verified from the above equation that ${\rm Tr}\left[H_{m}\mathcal{D}_{jk}(\rho)\right]=0$ for all $j,k$ except when $j=M$ and $k=m$. Therefore, Eq. (18) reduces to
${\rm Tr}\left[H_{m}\frac{d\rho}{dt}\right]$ $=-i{\rm Tr}\left[H_{m}H_{S}\rho-H_{m}\rho H_{S}\right]+{\rm Tr}\left[H_{m}\mathcal{D}_{Mm}(\rho)\right].$ (21)
Further, the first term on the RHS in the above equation can be rewritten as
$-i{\rm Tr}\left[H_{m}H_{S}\rho-H_{m}\rho H_{S}\right]$ $=-i{\rm Tr}\left[H_{m}H_{S}\rho-H_{S}H_{m}\rho\right]$
$=-i{\rm Tr}\left(\left[H_{m},H_{S}\right]\rho\right).$ (22)
The commutator of $H_{m}$ with $H_{S}$ can further be simplified to give $[H_{m},H_{S}]=-\sum_{l=1,l\neq m}^{n}[H_{lm},H_{m}]$, using which we can write
$-i{\rm Tr}\left[H_{m}H_{S}\rho-H_{m}\rho H_{S}\right]=\sum_{l=1,l\neq m}^{n}i{\rm Tr}\left(\left[H_{lm},H_{m}\right]\rho\right).$ (23)
Substituting the above equation in Eq. (21), we get the net current directed towards qubit $m$ as
${\rm Tr}\left[H_{m}\frac{d\rho}{dt}\right]=\sum_{l=1,l\neq m}^{n}\left\{i{\rm Tr}\left(\left[H_{lm},H_{m}\right]\rho\right)\right\}+{\rm Tr}\left[H_{m}\mathcal{D}_{Mm}(\rho)\right].$ (24)
It can be observed from the above equation that the heat current $\mathcal{J}_{lm}$ directed towards qubit $m$ from any qubit $l$ is given by
$\mathcal{J}_{lm}=i{\rm Tr}\left(\left[H_{lm},H_{m}\right]\rho\right),$ (25)
and the heat current $\mathcal{J}_{Mm}$ from bath $M$ (attached to qubit $m$) to qubit $m$ is given by
$\mathcal{J}_{Mm}={\rm Tr}\left[H_{m}\mathcal{D}_{Mm}(\rho)\right].$ (26)
Further, in the steady state condition, the net current directed towards qubit $m$, Eq. (24), becomes zero, that is, ${\rm Tr}\left[H_{m}\frac{d\rho^{SS}}{dt}\right]=0$. This implies,
$\sum_{l=1,l\neq m}^{n}\mathcal{J}_{lm}+\mathcal{J}_{Mm}=0.$ (27)
The above equation proves the quantum thermal version of Kirchhoff’s current law.
III Quantum Thermal Transformer
Here, we provide proof of the expression regarding the quantum thermal transformer. Consider the quantum thermal current between two qubits $\mathcal{J}_{jk}=i{\rm Tr}\left\{\rho^{SS}[H_{jk},H_{k}]\right\}$, where $H_{k}=\frac{\omega_{k}}{2}\sigma^{z}_{k}$ and $H_{jk}=J_{jk}\left(\sigma^{x}_{j}\sigma^{x}_{k}+\sigma^{y}_{j}\sigma^{y}_{k}\right),$ ($j\neq k$). Now,
$[H_{jk},H_{j}]$ $=H_{jk}H_{j}-H_{j}H_{jk}$
$=J_{jk}\frac{\omega_{j}}{2}\left(\sigma^{x}_{j}\sigma^{x}_{k}+\sigma^{y}_{j}\sigma^{y}_{k}\right)\sigma^{z}_{j}-J_{jk}\frac{\omega_{0}}{2}\sigma^{z}_{j}\left(\sigma^{x}_{j}\sigma^{x}_{k}+\sigma^{y}_{j}\sigma^{y}_{k}\right)$
$=J_{jk}\omega_{j}\left[-i\sigma^{y}_{j}\sigma^{x}_{k}+i\sigma^{z}_{j}\sigma^{x}_{k}\right],~{}~{}~{}~{}\text{and}$
$[H_{jk},H_{k}]$ $=H_{jk}H_{k}-H_{k}H_{jk}$
$=J_{jk}\frac{\omega_{k}}{2}\left(\sigma^{x}_{j}\sigma^{x}_{k}+\sigma^{y}_{j}\sigma^{y}_{k}\right)\sigma^{z}_{k}-J_{jk}\frac{\omega_{0}}{2}\sigma^{z}_{k}\left(\sigma^{x}_{j}\sigma^{x}_{k}+\sigma^{y}_{j}\sigma^{y}_{k}\right)$
$=J_{jk}\omega_{k}\left[i\sigma^{y}_{j}\sigma^{x}_{k}-i\sigma^{z}_{j}\sigma^{x}_{k}\right],$ (28)
From the above, we can verify that
$\frac{[H_{jk},H_{j}]}{\omega_{j}}=-\frac{[H_{jk},H_{k}]}{\omega_{k}}.$ (29)
Therefore, we get a relationship between the currents between two qubits $\mathcal{J}_{jk}$ (taking qubit $k$ as the node) and $\mathcal{J}_{kj}$ (taking qubit $j$ as the node) as
$\frac{\mathcal{J}_{jk}}{\omega_{k}}=-\frac{\mathcal{J}_{kj}}{\omega_{j}}.$ (30)
IV Thermal potentials
Consider a quantum thermal circuit made up of three qubits. The Hamiltonian for this three-qubit setup is given by
$\widetilde{H}_{S}$ $=H_{1}+H_{2}+H_{3}+H_{12}+H_{13}+H_{23},$
$=\frac{\omega_{1}}{2}\sigma_{1}^{z}+\frac{\omega_{2}}{2}\sigma_{2}^{z}++\frac{\omega_{3}}{2}\sigma_{3}^{z}+J_{12}\left(\sigma^{x}_{1}\sigma^{x}_{2}+\sigma^{y}_{1}\sigma^{y}_{2}\right)$
$+J_{13}\left(\sigma^{x}_{1}\sigma^{x}_{3}+\sigma^{y}_{1}\sigma^{y}_{3}\right)+J_{23}\left(\sigma^{x}_{2}\sigma^{x}_{3}+\sigma^{y}_{2}\sigma^{y}_{3}\right),$ (31)
where qubits $l$ and $k$ interact with each other via Heisenberg $XX$ type interaction with the interaction strength $J_{lk}$ and $\omega_{k}$ being the transition frequency of the $k$-th qubit. Under the Born-Markov and secular approximations, Eq. (11) with $\widetilde{H}_{S}$ in place of $H_{S}$ dictates the dynamics of the three-qubit system. By equating this to zero, we find the steady-state $\tilde{\rho}^{SS}$ of the three-qubit system. The effective temperatures of the qubits 1, 2, and 3 ($T_{1}$, $T_{2}$, and $T_{3}$, respectively), using the steady-state $\tilde{\rho}^{SS}$ of the three-qubit system are given by
$T_{1}$ $=\frac{\omega_{1}}{\log\left(\frac{1}{\tilde{\rho}^{SS}_{00}+\tilde{\rho}^{SS}_{11}+\tilde{\rho}^{SS}_{22}+\tilde{\rho}^{SS}_{33}}-1\right)},$ $T_{2}=\frac{\omega_{2}}{\log\left(\frac{1}{\tilde{\rho}^{SS}_{00}+\tilde{\rho}^{SS}_{11}+\tilde{\rho}^{SS}_{44}+\tilde{\rho}^{SS}_{55}}-1\right)},$
$T_{3}$ $=\frac{\omega_{3}}{\log\left(\frac{1}{\tilde{\rho}^{SS}_{00}+\tilde{\rho}^{SS}_{22}+\tilde{\rho}^{SS}_{44}+\tilde{\rho}^{SS}_{66}}-1\right)}.$ (32)
The corresponding thermal potential differences are given by
$V_{1A}$ $=T_{1}-0;$ $V_{21}=T_{2}-T_{1};$ $V_{32}=T_{3}-T_{2};$
$V_{13}$ $=T_{1}-T_{3};$ $V_{B2}=T-T_{2}.$ (33)
In terms of the elements of the steady state $\tilde{\rho}^{SS}$ of the three-qubit circuit, the quantum thermal currents between the qubits 1 and 3 and between qubits 1 and 2, taking qubit 1 as a node are given by
$\mathcal{J}_{31}$ $=-4J_{13}\omega_{1}\left[\Im{(\tilde{\rho}^{SS}_{14})}+\Im{(\tilde{\rho}^{SS}_{36})}\right],$
$\mathcal{J}_{21}$ $=-4J_{12}\omega_{1}\left[\Im{(\tilde{\rho}^{SS}_{24})}+\Im{(\tilde{\rho}^{SS}_{35})}\right],$ (34)
where $\Im(z)$ denotes imaginary part of $z$. Now, by defining a function
$\tilde{f}_{13}$ $=\frac{1}{\left[\Im{(\tilde{\rho}^{SS}_{14})}+\Im{(\tilde{\rho}^{SS}_{36})}\right]}\left[\frac{\omega_{3}}{\omega_{1}\log\left(\frac{1}{\tilde{\rho}^{SS}_{00}+\tilde{\rho}^{SS}_{22}+\tilde{\rho}^{SS}_{44}+\tilde{\rho}^{SS}_{66}}-1\right)}\right.$
$-\left.\frac{1}{\log\left(\frac{1}{\tilde{\rho}^{SS}_{00}+\tilde{\rho}^{SS}_{11}+\tilde{\rho}^{SS}_{22}+\tilde{\rho}^{SS}_{33}}-1\right)}\right],$ (35)
which is a function of $\omega_{i},T_{i},J_{ij},T,\gamma_{i}$ (for $i,j=1,2,3$), a relation between the quantum thermal heat current $\mathcal{J}_{31}$ and the thermal potential $V_{13}$ is given as
$\mathcal{J}_{31}=\frac{4J_{13}}{\tilde{f}_{13}}V_{13}.$ (36)
This is Eq. (7) in the main text. Further, using the relation $\frac{\mathcal{J}_{31}}{\omega_{1}}=-\frac{\mathcal{J}_{13}}{\omega_{3}}$, we can write
$\mathcal{J}_{13}=-\frac{4\omega_{3}J_{13}}{\omega_{1}\tilde{f}_{13}}V_{13}.$ (37)
Similarly, we can find out various functions, such as $\tilde{f}_{12}$, and $\tilde{f}_{23}$, which can be used to identify the corresponding relation of the quantum thermal currents $\mathcal{J}_{12},\mathcal{J}_{21},\mathcal{J}_{23}$, and $\mathcal{J}_{32}$ with the quantum thermal potentials $V_{12}$ and $V_{23}$.
V Quantum Thermal Diode
Here, we apply the framework developed to a few other quantum thermal circuits, for example, the quantum thermal diode [16, 53] followed by the quantum thermal transistor. The Hamiltonian of the quantum thermal diode is similar to the quantum thermal resistor discussed above, Eq. (10). The difference is that here, we take the temperature of bath $A$ to be non-zero. Accordingly, the dynamical master equation, Eq. (11), changes to accommodate the factor $N_{th,A}=\frac{1}{e^{\omega_{1}/T_{A}}-1}$ with the Lindblad operator $\sigma^{+}_{1}$. The steady-state for this system can be found similarly. We calculate the thermal heat current that flows between bath $A$ and the qubit 1. Based on the temperature difference between bath $A$ and bath $B$, we call the thermal diode to be in the reverse and in the forward bias. The voltage and, correspondingly, the thermal heat currents are zero when the temperatures of both $A$ and $B$ are the same.
Figure 7: Variation of heat current $\mathcal{J}_{A1}$ from bath $A$ to qubit 1 as a function of temperature $T_{A}$ of bath $A$ for a quantum thermal diode. The parameters are: $\omega_{1}=1.5,\omega_{2}=1.5,J_{12}=0.5,\gamma_{A}=0.01,\gamma_{B}=0.1,T_{B}=1.0$.
In Fig. 7, we show the variation of quantum thermal heat current $\mathcal{J}_{A1}$ with temperature of bath $A$. The forward and the reverse bias regions are specified explicitly in the figure. In the forward bias region, the heat current flows from bath $A$ to the qubit 1, whereas in the reverse bias region, the heat current flows from the qubit 1 to the bath $A$. Further, the heat current is zero when the temperatures $T_{A}$ and $T_{B}$ of the baths $A$ and $B$, respectively, are equal.
VI Quantum Thermal Transistor
The quantum thermal transistor circuit is an upgrade to the quantum thermal diode [17, 18]. It also acts as an amplifier of the heat current. The circuit model requires three qubits and three baths.
Figure 8: A schematic diagram to show the arrangement of qubits and the respective baths to function as a quantum thermal transistor. The labels $B$, $C$, and $E$ correspond to base, collector, and emitter, respectively.
A schematic diagram of the circuit is provided in Fig. 8. The labels $B$, $C$, and $E$ correspond to base, collector, and emitter, respectively.
Figure 9: Variation of the (a) collector ($\mathcal{J}_{C}$), base ($\mathcal{J}_{B}$), and emitter ($\mathcal{J}_{E}$) thermal currents and (b) the corresponding amplification factor $\alpha_{C}$ and $\alpha_{E}$ as a function of the base temperature $T_{B}$. The parameters take the following values: $\omega_{1}=\omega_{2}=1.0,\omega_{3}=0.05\omega_{1},J_{12}=J_{13}=J_{23}=1.0,T_{E}=T_{C}=0.2,\gamma_{E}=0.003,\gamma_{B}=0.01,\gamma_{C}=0.002$.
Further, the three baths have non-zero temperatures. The interaction strength factors between the qubits $B$, $C$, and $E$ are $J_{BC}$, $J_{EB}$, and $J_{CE}$. The Hamiltonian of the system is given by
$H_{trans}$ $=\frac{\omega_{C}}{2}\sigma^{z}_{C}+\frac{\omega_{B}}{2}\sigma^{z}_{B}+\frac{\omega_{E}}{2}\sigma^{z}_{E}+J_{BC}\left(\sigma^{x}_{B}\sigma^{x}_{C}+\sigma^{y}_{B}\sigma^{y}_{C}\right)$
$+J_{EB}\left(\sigma^{x}_{E}\sigma^{x}_{B}+\sigma^{y}_{E}\sigma^{y}_{B}\right)+J_{CE}\left(\sigma^{x}_{C}\sigma^{x}_{E}+\sigma^{y}_{C}\sigma^{y}_{E}\right).$ (38)
The above Hamiltonian is along similar lines to our adaptation of thermal resistors, diodes, and other thermal circuits discussed here. The master equation dictating the dynamics of the system is given by
$\frac{d\rho}{dt}$ $=-i[H_{trans},\rho]+\sum_{k=B,C,E}\left(\gamma_{k}\left(N_{th,k}+1\right)\left[\sigma^{-}_{k}\rho\sigma^{+}_{k}-\frac{1}{2}\left\{\sigma^{+}_{k}\sigma^{-}_{k}\right\}\right]\right.$
$+\left.\gamma_{k}N_{th,k}\left[\sigma^{+}_{k}\rho\sigma^{-}_{k}-\frac{1}{2}\left\{\sigma^{-}_{k}\sigma^{+}_{k}\right\}\right]\right),$ (39)
where $N_{th,k}=\frac{1}{e^{\omega_{k}/T_{k}}-1}$ for $k=B,C,E$. The base, emitter, and collector currents $\mathcal{J}_{B}$, $\mathcal{J}_{E}$, and $\mathcal{J}_{C}$, respectively, for the quantum thermal transistor are found using the steady-state $\rho^{SS}_{trans}$ in Eq. (26) and are plotted in Fig. 9(a) as a function of the base temperature $T_{B}$. It can be observed that the base current only changes slightly, whereas, for the same change in the temperature, the collector and the emitter currents change rapidly. To quantify the amplifications in the collector and emitter currents, the factor
$\alpha_{P}=\frac{\mathcal{J}_{P}}{\mathcal{J}_{B}},$ (40)
for $P=C,E$ is plotted in Fig. 9(b). It can be observed that the collector and the emitter currents get amplifications of 8 and 12 times, respectively, for the given set of parameters.